OpenStudy (anonymous):

Whitney left school and drove 7 miles to meet her mother at a restaurant. Her mother left at the same time and drove 8 miles, traveling 5 mi/h faster than her daughter. If the two arrived at the restaurant at the same time, what was Whitney’s speed? a 30 mi/h b 35 mi/h c 46 mi/h d 51 m/h

OpenStudy (anonymous):

@johnweldon1993 can u help?

OpenStudy (zpupster):

if you could not solve what i gave i would of helped

OpenStudy (anonymous):

sorry

OpenStudy (johnweldon1993):

We know that Distance = rate times time \[\large D = rt\] or we can write this in terms of time (you'll see why in a few) \[\large t = \frac{d}{r}\] The time Whitney traveled we'll write as \[\large t_w = \frac{d_w}{r_w}\] and the time for her mother lets write as \[\large t_m = \frac{d_m}{r_m}\] We also know that the travel took the same time for both...so So what we now have is \[\large t_w = t_m\] lets plug in what we know..and that is \[\large \frac{d_w}{r_w} = \frac{d_m}{r_m}\] so...we know Whiteny trveled 7 miles....and the other traveled 8 \[\large \frac{7}{r_w} = \frac{8}{r_m}\] lets say whitney traveled 'x' mph and her mother would have done x + 5 (since she traveled 5mph faster than her daughter) so \[\large \frac{7}{x} = \frac{8}{x + 5}\] Now we just cross multiply and solve for 'x'

OpenStudy (anonymous):

ok

OpenStudy (johnweldon1993):

Let me know what you get :)

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

35?

OpenStudy (johnweldon1993):

Perfect!

OpenStudy (anonymous):

so thats the answer

OpenStudy (johnweldon1993):

indeed^ :)

OpenStudy (anonymous):

thanks XD

OpenStudy (johnweldon1993):

lol no problem :P